# Jumpstarting Surgical Computer Vision

### v1

@@ 4 Conclusion
Recent developments in self-supervised learning for surgery allow us to learn generic representational features, useful for a range of downstream tasks in the domain. In this work, we build on these developments, shifting the focus from models to data, and exploring whether pre-trained networks could be an effective means to consolidate disparate datasets in the field of Surgical Data Science. We illustrate that simple model initializations can indeed help leverage these disparate datasets, helping ’jumpstart’ work on new problems in new procedures at new clinical centers with fewer annotations.
5 Acknowledgments
This work was supported by French state funds managed by the ANR within the National AI Chair program under Grant ANR-20-CHIA-0029-01 (Chair AI4ORSafety) and within the Investments for the future program under Grant ANR-10-IAHU-02 (IHU Strasbourg), and by BPI France under reference DOS0180017/00 (project 5G-OR). It was granted access to the HPC resources of Unistra Mesocentre. This work was granted access to the HPC resources of IDRIS under the allocations 2022-AD011011640R2, 2023-AD011013523R1, and 2023-AD011011631R3 made by GENCI. The AI4SafeChole consortium is represented by Giovanni Guglielmo Laracca, Ludovica Guerriero, Andrea Spota, Claudio Fiorillo, Giuseppe Quero, Segio Alfieri, Ludovica Baldari, Elisa Cassinotti, Luigi Boni, Diego Cuccurullo, Guido Costamagna, and Bernard Dallemagne.

### v2

@@ 4 Conclusion
Recent developments in self-supervised learning for surgery allow us to learn generic representational features, useful for a range of downstream tasks in the domain. In this work, we build on these developments, shifting the focus from models to data, and explore whether pre-trained networks could be an effective means to consolidate disparate datasets in the field of Surgical Data Science. We try to illustrate, that simple model initializations could help promote work on new problems in new procedures at new clinical centers with fewer annotations. The pre-trained initializations generated through this work will be made available alongside the released code at https://github.com/CAMMA-public/ScalingSurgicalSSL/.
5 Acknowledgement
This work was supported by French state funds managed by the ANR under Grant ANR-20-CHIA-0029-01 (Chair AI4ORSafety) and Grant ANR-10-IAHU-02 (IHU Strasbourg), and by BPI France under reference DOS0180017/00 (project 5G-OR). The AI4SafeChole consortium is represented by Ludovica Baldari, Giovanni Guglielmo Laracca, Ludovica Guerriero, Segio Alfieri, Claudio Fiorillo, Giuseppe Quero, Elisa Cassinotti, Luigi Boni, Diego Cuccurullo, Guido Costamagna, and Bernard Dallemagne.
[…]
The authors have no competing interests to declare.